High-Low Method for a Service Company Continental Railroad decided to use the high-low method and operating...

  1. High-Low Method for a Service Company

    Continental Railroad decided to use the high-low method and operating data from the past six months to estimate the fixed and variable components of transportation costs. The activity base used by Continental Railroad is a measure of railroad operating activity, termed “gross-ton miles,” which is the total number of tons multiplied by the miles moved.

    Transportation Costs Gross-Ton Miles
    January $24,500,000 3,000,000
    February 22,375,000 2,500,000
    March 29,000,000 6,300,000
    April 34,800,000 9,500,000
    May 40,312,500 12,750,000
    June 35,500,000 10,000,000

    Determine the variable cost per gross-ton mile and the total fixed cost.

    Variable cost (If required, round to two decimal places.) $ per gross-ton mile
    Total fixed cost $
Answer #1

Variable cost per gross-ton mile = (Highest activity cost - Lowest activity cost)/(Highest activity - Lowest activity)

= (40,312,500 - 22,375,000)/(12,750,000 - 2,500,000)

= 17,937,500/10,250,000

= $1.75 per gross-ton mile

Total Fixed cost = Highest activity cost - (Highest activity x Variable cost per gross-ton mile)

= 40,312,500 - (12,750,000 x 1.75)

= 40,312,500 - 22,312,500

= $18,000,000